# FedFitTech: A Baseline in Federated Learning for Fitness Tracking

### v1

@@ 3.2. Dataset
We used inertial-based data of the WEAR (Wearable and Egocentric Activity Recognition) (Bock et al., 2024) dataset which consists of labelled activities performed by 22 participants and categorized into three types: jogging, stretching, and strength exercises. The jogging category includes five labels: normal, rotating arms, skipping, sidesteps, and butt-kicks. The stretching category also comprises five labels: triceps, lunging, shoulders, hamstrings, and lumbar rotation. Lastly, the strength category encompasses eight labels: push-ups, push-ups (complex), sit-ups, sit-ups (complex), burpees, lunges, lunges (complex), and bench dips. Also, the dataset includes a NULL label that represents unlabeled or undefined periods, where no activity was explicitly carried out. It serves as a placeholder for missing annotations and can be useful for filtering, handling uncertainty, or distinguishing between known and unknown states. As a result, participants engaged in 18 different outdoor sports activities. Unlike previous egocentric datasets, WEAR features an inertial sensor placement aligned with recent real-world application trends. It is the first dataset collected across multiple outdoor locations, each presenting varying surface conditions without revealing cues about the performed activities. Moreover, in FL in HAR applications, the selected dataset must mirrors real-world conditions, as it should accommodate a large number of clients and a diverse array of classes and subjects. Herewith, based on the recent work (Geissler et al., 2024), the WEAR dataset remains the best option for fitness tracking since other datasets cannot fulfil needs due to having fewer participants, classes and labelled data.
@@ 3.3. Hyperparameters and Framework
Client setting: In the dataset, subject 1 and 19, and also subject 15 and 20 are from the same participants. These two participants were re-recorded in a different season and environment, resulting in 24 subjects in total. These two additional subjects are treated as separate devices for the same individuals to reflect real-world FitTech scenarios, thus setting the number of clients to 24.
Data splitting: In the WEAR dataset, participants perform the same activity several times in different time periods randomly, to represent realistic scenarios. This means that the activity patterns of people can change over time. For instance, while a person jogs fast at the beginning of exercise, he/she can feel tired later which causes body movements to be slower. Furthermore, considering factors such as weather conditions and changes on the ground, it cannot be expected that a person will make the same body movements. Therefore, the method of splitting local data into training and testing requires special scrutiny. For this reason, we select the first 20 % of each label in the time series to add to testing data and the remaining 80 % is left for training. In this way, for instance, when a person jogs at the beginning of exercise and also at a later time of exercise, testing data will include these two patterns.
Window size: After conducting multiple experiments, we determined that a window size of 100 is optimal option, corresponding to a duration of 2 seconds in the time series sampled at 50 Hz.
Batch size: Despite the impact of other hyperparameters differing based on the utilized model and dataset, (Kundroo and Kim, 2024) reported that a batch size of 32 emerged as the optimal choice for image data. Nevertheless, we set to 32, as the hyperparameter recommendation for time series remains underexplored, and this value used in literature as well (Tu et al., 2021).
Optimizer: We employed the Adam optimizer because it has shown to converge quickly and efficiently for complex models and large datasets. By adaptively tuning each parameter’s learning rate and using momentum to smooth the optimization, it requires less hyperparameter tuning and these were found to make it well-suited for Federated Learning contexts (Li et al., 2021; Presotto et al., 2022).
@@
Global round: Based on existing work in Federated Learning in HAR (Ek et al., 2022), we have set the global round to 100, as it is also a standard setting for other domains in FL in general.
@@
Early stopping: Inspired by (Prechelt, 2002), we implemented early stopping based on the stability of the validation F1-score over a sliding window. The hyperparameters of this method are set based on our several experiments, to determine the optimum values for this paper’s case study. As a result of experimentation, the patience value is set to 5, and the threshold is set to 0.01 for the stopping criterion.
@@ 4. Evaluation Results
Evaluation of this work compares FedFitTech with its client-side early stopping case study, considering both in communication loads and global model performance on local data.
@@ 4.1. Communication Cost
Fig. 2 illustrates that 9 out of 24 of clients stopped early from training global model, highlighting that prolonged participating does not guarantee continued benefit. The earliest stop was round 40 and some clients decided to stop early even though they had participated in more than 80 training rounds.
Figure 2. Amount of training rounds each client attended in the case study. The Y-axis shows the F1-score and the X-axis shows the client IDs. Red bars depicts the communication cost, while green bars show the saved communication rounds, and the dashed black line is the mean for these.
Figure 3. Communication rounds (X-axis) versus F1-scores (Y-axis). The dashed black line shows the mean values of all local performance, the triangles markers depict early stopped clients rounds.
Moreover, Fig. 3 shows F1-score performance of clients which are dropped early (see triangle markers within the plot). Some clients could not increase their local F1-score even for 30 rounds of training. Also, as can be seen from the black dashed line, the overall global model F1-score continues to increase, despite some clients being dropped.
@@ 4.2. Model Performance
The client-based F1-scores are depicted in Fig. 4 for both the baseline and our case study of early stopping. The FedFitTech baseline resulted in a 68% mean F1-score across all clients, with the early stopping case study performing closely with a mean F1-score of 67%. Additionally, the result presents that some of the clients even have better F1-scores in the system in the case study, such as the clients represented with these ID numbers: 2, 7, 9, 10, 11, 12, 14, 17, 19, 21, and 22, with 11 out of 24 clients obtaining a better performance compared to the FedFitTech baseline.
Besides, Fig. 5 and Fig. 6 depict the label based F1-scores of FedFitTech baseline and the case study, respectively. Actual label names in these figures are as follows: ’A’: ’NULL’, ’B’: ’jogging’, ’C’: ’jogging (rotating arms)’, ’D’: ’jogging (skipping)’, ’E’: ’jogging (sidesteps)’, ’F’: ’jogging (butt-kicks)’, ’G’: ’stretching (triceps)’, ’H’: ’stretching (lunging)’, ’I’: ’stretching (shoulders)’, ’J’: ’stretching (hamstrings)’, ’K’: ’stretching (lumbar rotation)’, ’L’: ’push-ups’, ’M’: ’push-ups (complex)’, ’N’: ’sit-ups’, ’O’: ’sit-ups (complex)’, ’P’: ’burpees’, ’Q’: ’lunges’, ’R’: ’lunges (complex)’, and ’S’: ’bench-dips’.
@@
Figure 6. The client-label F1-scores for this paper’s case study of early stopping. When comparing with Fig. 5, it can be seen that overall classification performance remains similar, with certain activities (such as L - push ups and M - complex push ups) remaining challenging across users.
Despite early stopping decreasing the overall mean F1-score slightly by 1% (see Fig. 4), some of the clients reached a remarkable performance rise for some of the labels. For instance, the push-up fitness activity (see label L in the confusion matrices) had an F1-score for Client 2 in FedFitTech of 0%, while it is 76% in the case study. Client 18, which was the one that stopped training the earliest (see Fig. 2), achieved a 47% reduction in communication costs and its overall F1-score increased from 71% to 74%. Also, tracking some of the activities (e.g. ’stretching (lunging)’: H) increased from 55% to 86%.
@@ 4.3. Discussion
Fig. 2 and 3 show that this client attrition, while seemingly modest in this small-scale experiment, represents a significant concern in real-world Federated Learning deployments involving a large number of clients, as 37.5% of clients decreased their communication loads. Besides, these findings suggest that client-specific factors, such as data heterogeneity, resource constraints, or local model convergence, may influence participation stability within the Federated Learning framework. However, as can be seen from Fig. 3, the global model continues to improve despite some clients not longer participating in training. Also, Fig. 4 demonstrates that in the case study, 45.8% of clients increased their overall mean F1-scores while communication costs decreased by 13% (see Fig. 2).
@@ 5. Conclusions and Future Work
Wearable devices provide rich data that enable in-depth insights into users’ fitness activities within the domain of Fitness Technology (FitTech). However, manual data annotation pose significant challenges for FitTech. For this reason, the user could benefit from other people annotating activities on their wearable clients. Centralized Learning, while powerful, faces issues such as latency, scalability, and privacy.
We argue in this paper that Federated Learning fits well with the characteristics of FitTech, due to clients having largely similar sensors, similar processing power, and mostly common activity classes, while needing relatively infrequent communication between a global model and the clients. This work thus introduces FedFitTech, designed to explore the FitTech domain within Federated Learning, built upon the user-friendly, widely adopted, and open-source Flower framework. Moreover, we shared a case study as an example usage of this baseline which involves a client-side early stopping strategy. The case study is compared with the FedFitTech baseline, considering communication loads and model performance metrics. The finding shows that our case study mitigates communication burden with a negligible mean F1-score drop, by balancing model generalization.
The FedFitTech baseline provides various promising directions for future research in Federated Learning for the FitTech domain. For instance, exploring transfer learning implementations can accelerate convergence and enable clients to benefit from pre-existing patterns from the outset. Additionally, integrating differential privacy techniques is crucial for enhancing data privacy and security, addressing a fundamental concern in FitTech. Furthermore, since the utilized WEAR dataset in FedFitTech consists of both inertial and ego-perspective camera data, investigating multimodal mechanisms has significant potentials for handling diverse sensor data and mitigating data heterogeneity. In conclusion, we hope that our FedFitTech baseline will leverage the publicly available and widely used Flower framework, opening up a new research domain and development opportunities in FitTech.

### v2

@@ 3.2. Dataset
We used inertial-based data from the WEAR (Wearable and Egocentric Activity Recognition) (https://mariusbock.github.io/wear/, (Bock et al., 2024)) dataset, which consists of labeled activities performed by 22 participants and categorized into three types: jogging, stretching, and strength exercises. The jogging category includes five labels: normal, rotating arms, skipping, sidesteps, and butt-kicks. The stretching category also comprises five labels: triceps, lunging, shoulders, hamstrings, and lumbar rotation. Lastly, the strength category encompasses eight labels: push-ups, push-ups (complex), sit-ups, sit-ups (complex), burpees, lunges, lunges (complex), and bench dips. In addition, the dataset includes a NULL label that represents unlabeled or undefined periods where no activity was explicitly carried out. It serves as a placeholder for missing annotations and can be useful for filtering, handling uncertainty, or distinguishing between known and unknown states. As a result, participants engaged in 18 different outdoor sports activities. Unlike previous egocentric datasets, WEAR features an inertial sensor placement aligned with recent real-world application trends. It is the first dataset collected across multiple outdoor locations, each presenting varying surface conditions without revealing cues about the performed activities. Moreover, in FL in HAR applications, the selected dataset must mirror real-world conditions, as it should accommodate a large number of clients and a diverse array of classes and subjects. Herewith, based on the recent work (Geissler et al., 2024), the WEAR dataset remains the best option for fitness tracking, since other datasets cannot fulfill the needs due to having fewer participants, classes, and labeled data.
@@ 3.3. Hyperparameters and Framework
Client setting: In the dataset, subjects 1 and 19, and also subjects 15 and 20, are from the same participants. These two participants were re-recorded in a different season and environment, resulting in 24 subjects in total. These two additional subjects are treated as separate devices for the same individuals to reflect real-world FitTech scenarios, thus setting the number of clients to 24.
Data splitting: In the WEAR dataset, participants perform the same activity several times in different time periods randomly, to represent realistic scenarios. This means that the activity patterns of people can change over time. For instance, while a person jogs fast at the beginning of exercise, he/she can feel tired later, which causes body movements to be slower. Furthermore, considering factors such as weather conditions and changes on the ground, it cannot be expected that a person will make the same body movements. Therefore, the method of splitting local data into training and testing requires special scrutiny. For this reason, we select the first 20 % of each label in the time series to add to the testing data, and the remaining 80 % is left for training. In this way, for example, when a person jogs both at the start and later in the session, the test set will include both jogging bouts, ensuring temporal variation within each class.
Window size: After conducting multiple experiments, we determined that a window size of 100 is the optimal option, corresponding to a duration of 2 seconds in the time series sampled at 50 Hz.
Batch size: Despite the impact of other hyperparameters differing based on the used model and the dataset, (Kundroo and Kim, 2024) reported that a batch size of 32 emerged as the optimal choice for image data. Nevertheless, we set it to 32, since the hyperparameter recommendation for time series remains underexplored, and this value is used in the literature as well (Tu et al., 2021).
Optimizer: We employed the Adam (Adaptive Moment Estimation) optimizer because it has been shown to converge quickly and efficiently for complex models and large datasets. By adaptively tuning the learning rate of each parameter and using momentum to smooth the optimization, it requires less hyperparameter tuning, and these were found to be suitable for Federated Learning contexts (Li et al., 2021; Presotto et al., 2022).
@@
Communication round: Based on the existing work in Federated Learning in HAR (Ek et al., 2022), we have set the communication round to 100, as it is also a standard setting for other domains in FL in general.
@@
Early stopping: Following the method in (Prechelt, 2002), we implemented early stopping based on the stability of the validation F1-score over a sliding window. The hyperparameters of this method are set based on our several experiments to determine the optimum values for this paper’s case study. As a result of experimentation, the patience value is set to 5, and the threshold is set to 0.01 for the stopping criterion.
@@ 4. Evaluation Results
The evaluation of this work compares FedFitTech with its client-side early stopping case study, considering both communication loads and global model performance on local data.
Figure 2. Amount of training rounds each client attended in the case study. The Y-axis shows the communication round, and the X-axis shows the client IDs. Red bars depict the communication cost, while green bars show the saved communication rounds, and the dashed black line is their mean.
@@ 4.1. Communication Cost
Fig. 2 illustrates that 9 out of 24 clients stopped early from training the global model, highlighting that prolonged participation does not guarantee continued benefit. The earliest stop was observed at the 40th communication round (see client ID 3 in Fig. 2), and some clients decided to stop early, even though they had participated in training the global model in more than 80 communication rounds.
Figure 3. Communication rounds (X-axis) versus F1-scores (Y-axis). The dashed black line shows the mean values of all local performance, and the triangle markers depict rounds of early stopped clients.
Moreover, Fig. 3 shows the performance of the F1-score of clients who are dropped early (see triangle markers within the plot). Some clients were unable to increase their local F1-score even after 30 rounds of training. Also, as can be seen from the black dashed line, the overall global model F1-score continues to increase, despite some clients being dropped.
@@ 4.2. Model Performance
The client-based F1-scores are depicted in Fig. 4 for both the baseline and our case study of early stopping. The FedFitTech baseline resulted in a 68% mean F1-score across all clients, with the early stopping case study performing closely with a mean F1-score of 67%. Additionally, the result presents that some of the clients even have better F1-scores in the system in the case study, such as the clients represented with these ID numbers: 2, 7, 9, 10, 11, 12, 14, 17, 19, 21, and 22, with 11 out of 24 clients obtaining a better performance compared to the FedFitTech baseline. In conclusion, the overall recognition performance remains almost unchanged, while early stopping reduces communication costs and overall computational load.
@@
Figure 6. The client-label F1-scores for this paper’s case study of early stopping. When comparing with Fig. 5, it can be seen that overall classification performance remains similar, with certain activities (such as L - push-ups and M - complex push-ups) remaining challenging across users.
Besides, Fig. 5 and Fig. 6 depict the label-based F1 scores of the FedFitTech baseline and the case study, respectively. Actual label names in these figures are as follows: ’A’: ’NULL’, ’B’: ’jogging’, ’C’: ’jogging (rotating arms)’, ’D’: ’jogging (skipping)’, ’E’: ’jogging (sidesteps)’, ’F’: ’jogging (butt-kicks)’, ’G’: ’stretching (triceps)’, ’H’: ’stretching (lunging)’, ’I’: ’stretching (shoulders)’, ’J’: ’stretching (hamstrings)’, ’K’: ’stretching (lumbar rotation)’, ’L’: ’push-ups’, ’M’: ’push-ups (complex)’, ’N’: ’sit-ups’, ’O’: ’sit-ups (complex)’, ’P’: ’burpees’, ’Q’: ’lunges’, ’R’: ’lunges (complex)’, and ’S’: ’bench-dips’.
Despite early stopping decreasing the overall mean F1-score slightly by 1% (see Fig. 4), some of the clients reached a remarkable performance increase for some of the labels. For instance, the push-up fitness activity (see label L in the confusion matrices) had an F1-score for Client 2 in FedFitTech of 0%, while it is 76% in the case study. Client 18, which was the one who stopped training the first (see Fig. 2), achieved a 47% reduction in communication costs, and its overall F1-score increased from 71% to 74%. Also, tracking some of the activities (e.g., ’stretching (lunging)’: H) increased from 55% to 86%.
@@ 4.3. Discussion
Fig. 2 and 3 show that this client attrition, while seemingly modest in this small-scale experiment, represents a significant concern in real-world Federated Learning deployments involving a large number of clients, as 37.5% of clients decreased their communication loads. Besides, these findings suggest that client-specific factors, such as data heterogeneity, resource constraints, or local model convergence, may impact the stability of participation within the Federated Learning framework. However, as shown in Fig. 3, the global model continues to improve despite some clients no longer participating in training. Also, Fig. 4 demonstrates that in the case study, 45.8% of the clients increased their overall mean F1-scores, while communication costs decreased by 13% (see Fig. 2).
@@ 5. Conclusions and Future Work
Wearable devices provide rich data that enable in-depth insights into user fitness activities within the domain of Fitness Technology (FitTech). However, manual data annotation in FitTech is both time‑consuming and prone to error, creating a significant bottleneck. Users can offload the tedious task of annotating activities while benefiting from diverse and high quality labels. Although Centralized Learning approaches can harness all this annotated data, they often struggle with latency, limited scalability, and heightened privacy risks.
We argue in this paper that Federated Learning aligns well with the characteristics of FitTech, as clients have largely similar sensors, similar processing power, and mostly common activity classes, while requiring relatively infrequent communication between a global model and the clients. This work thus introduces FedFitTech, designed to explore the FitTech domain within Federated Learning, built upon the user-friendly, widely adopted, and open-source Flower framework. Moreover, we shared a case study as an example usage of this baseline, which involves a client-side early stopping strategy. The case study is compared with the FedFitTech baseline, considering communication loads and model performance metrics. The findings show that our case study mitigates the communication burden with a negligible mean F1-score drop by balancing model generalization.
The FedFitTech baseline provides various promising directions for future research in Federated Learning for the FitTech domain. For instance, exploring transfer learning implementations can accelerate convergence and enable clients to benefit from pre-existing patterns from the outset. Additionally, integrating differential privacy techniques is crucial for improving data privacy and security, addressing a fundamental concern in FitTech. Furthermore, since the utilized WEAR dataset in FedFitTech consists of both inertial and ego-perspective camera data, investigating multimodal mechanisms has significant potential for handling diverse sensor data and mitigating data heterogeneity.
Moreover, exploring efficient local model clustering methods and then sharing cluster-specific global models with clients could reduce performance differences across clients and increase overall recognition performance by facilitating collaboration among users with similar behavioral patterns (e.g., activity routines, age, weight, or gender). In conclusion, we hope that our FedFitTech baseline will leverage the publicly available and widely used Flower framework, opening up a new research domain and development opportunities in FitTech.
The FedFitTech baseline is available as an open source repository at: https://github.com/shreyaskorde16/FedFitTech